# Exposing the Cracks: Vulnerabilities of Retrieval-Augmented LLM-based Machine Translation

### v1

@@ Abstract
REtrieval-Augmented LLM-based Machine Translation (REAL-MT) shows promise for knowledge-intensive tasks like idiomatic translation, but its reliability under noisy retrieval contexts remains poorly understood despite this being a common challenge in real-world deployment. To address this gap, we propose a noise synthesis framework and new metrics to evaluate the robustness of REAL-MT systematically. Using this framework, we instantiate REAL-MT with Qwen-series models, including standard LLMs and large reasoning models (LRMs) with enhanced reasoning, and evaluate their performance on idiomatic translation across high-, medium-, and low-resource language pairs under synthesized noise. Our results show that low-resource language pairs, which rely more heavily on retrieved context, degrade more severely under noise than high-resource ones and often produce nonsensical translations. Although LRMs possess enhanced reasoning capabilities, they show no improvement in error correction and are even more susceptible to noise, tending to rationalize incorrect contexts. We find that this stems from an attention shift away from the source idiom to noisy content, while confidence increases despite declining accuracy, indicating poor calibration. To mitigate these issues, we investigate training-free and fine-tuning strategies, which improve robustness at the cost of performance in clean contexts, revealing a fundamental trade-off. Our findings highlight the limitations of current approaches, underscoring the need for self-verifying integration mechanisms.
@@ Introduction
REtrieval-Augmented LLM-based Machine Translation (REAL-MT) is increasingly used to enhance translation quality for knowledge-intensive MT tasks like idiomatic translation (Li et al. 2024; Donthi et al. 2025). Although external knowledge can enhance translation performance, reliance on it is a double-edged sword: when the retrieved context contains noise such as irrelevant or misleading information, LLMs often produce nonsensical translations, as illustrated in Figure 1. Since noisy retrieval is unavoidable in real-world deployment, the behavior of REAL-MT under such conditions remains poorly understood, posing a critical barrier to deploying REAL-MT in safety-sensitive or real-world applications where reliability is paramount. This gap motivates our central question: to what extent does noisy retrieval compromise REAL-MT’s trustworthiness?
@@
To investigate how noisy retrieval compromises REAL-MT’s trustworthiness, we develop a two-pronged approach: a controlled noise injection framework to simulate realistic retrieval failures, and specialized evaluation metrics to quantify their impact. While standard machine translation metrics like COMET (Rei et al. 2020) assess overall output quality, they fail to capture semantic fidelity in idiomatic translation and cannot distinguish whether errors stem from source misinterpretation or over-reliance on retrieved context. To address this gap, we propose two complementary metrics: Fidelity, which evaluates translation correctness with a focus on idiomatic accuracy, and Context Adoption Rate (CAR), which quantifies the extent to which models rely on external context. Together, they enable fine-grained analysis of both what the model gets wrong and why.
@@
We evaluate REAL-MT systems instantiated with standard LLMs and large reasoning models (LRMs) across high-, medium-, and low-resource language pairs under controlled synthesized noise. Our results show that REAL-MT is far more vulnerable to knowledge-level errors, such as irrelevant or contradictory retrieval content, than to surface-level perturbations (e.g., word reordering). Moreover, the extent of performance degradation scales with the degree of semantic deviation: the more the retrieved context diverges from the intended meaning of idioms, the more severe the drop in translation fidelity. This sensitivity is especially pronounced in low-resource language pairs, as reflected in their higher CAR, which signal greater dependence on external context and leads to more severe degradation under noisy retrieval. Surprisingly, LRMs, despite their enhanced reasoning capabilities, show no improvement in error correction and are even more susceptible to noise, often rationalizing incorrect contexts. Across noise conditions, CAR remains consistently high even when retrieved content contradicts the source, prompting us to examine whether this reflects active model integration or coincidental similarity. Our attention analysis confirms the former: models consistently attend to retrieved context regardless of its correctness, demonstrating active integration. This uncritical reliance is further exacerbated by poor metacognitive awareness, as confidence increases despite declining accuracy, a sign of severe miscalibration and absent self-verification.
@@
Given this uncritical reliance on retrieved context, we explore both training-free and fine-tuning strategies to improve REAL-MT robustness. While both enhance noise resistance, they incur a consistent trade-off: performance degrades under clean, accurate contexts. Fine-tuning yields better robustness overall but still fails to adjust reliance based on context quality. This persistent trade-off reveals that post-hoc mitigation cannot overcome the model’s fundamental inability to self-verify, underscoring the need for self-verifying integration mechanisms that validate retrieved content before adoption and enable rejection of noise while preserving accurate knowledge.
@@
In summary, this work (1) introduces a controlled noise injection framework to evaluate REAL-MT robustness under realistic retrieval failures systematically, (2) proposes Fidelity and Context Adoption Rate (CAR) as fine-grained metrics to diagnose error sources in knowledge-intensive translation, and (3) evaluates training-free and fine-tuning mitigation strategies, revealing a fundamental trade-off between robustness under noise and performance in clean contexts due to uncritical reliance on retrieval and poor calibration, which underscores the need for self-verifying integration mechanisms in REAL-MT systems.
@@
Large language models (LLMs) have revolutionized machine translation (MT), especially for low-resource languages or domains where sufficient parallel corpora are lacking. However, when confronted with translation scenarios demanding specific background knowledge, relying solely on the LLM’s internal knowledge proves inadequate. In such instances, incorporating external knowledge to address the inherent limitations of LLMs becomes crucial (Merx et al. 2024; Chen et al. 2024; Zebaze, Sagot, and Bawden 2025). The prompt engineering capabilities of LLMs enable the integration of externally retrieved knowledge via prompting, without additional training. This makes prompt-based retrieval-augmented generation an efficient and flexible solution for MT. Recent studies by Li et al. (2024) and Donthi et al. (2024) leverage LLMs to enhance idiomatic translation by incorporating idiom-meaning pairs retrieved from offline knowledge bases directly into the prompt.
@@
Retrieval-Augmented Language Models (RALM), reliant on external retrieved content, are susceptible to compromised reliability and robustness in their generated outputs when exposed to noise, irrelevant information, or malicious data  (Zhou et al. 2024; Park and Lee 2024; Shen et al. 2024; Yang et al. 2025). In response to the robustness challenges posed by RALMs, researchers have developed diverse approaches to enhance system robustness. Fang et al. (2024) introduces a named Retrieval-Augmented Adaptive Adversarial Training (RAAT) method to enhance the model’s ability to recognize and handle various types of noise.  Yoran et al. (2023) fine-tune the model using the parameter-efficient fine-tuning method QLoRA (Quantized Long-Range Attention) by using the synthetically generated noisy data to enhance its robustness to noisy data.  Xia et al. (2025) introduce a novel end-to-end self-reasoning framework that enhances the robustness, interpretability, and traceability of RALMs. This improvement is achieved by leveraging the reasoning trajectories generated by the LLMs themselves.
@@
To analyze how resource availability affects the robustness of retrieval-augmented LLM-based machine translation (MT), we group translation directions into high-, medium-, and low-resource tiers following Joshi et al. (2020), based on parallel data availability and typological distance from English. We compile a dataset of idiomatic translations spanning ten language pairs, selected only from sources that provide reference translations or explicit semantic annotations, ensuring reliable interpretation of idioms and enabling controlled noise synthesis. The collection integrates multiple publicly available resources:
  • High-resource: IdiomsInCtx-MT (Stap et al. 2024) (English–German, German–English, Russian–English), the French, Japanese, and Korean idioms from Liu, Chaudhary, and Neubig (2023) and the KISS dataset [^1] https://github.com/Judy-Choi/KISS-Korean-english-Idioms-in-Sentences-dataSet, all paired with English;
  • Medium-resource: the Finnish–English idioms from Liu, Chaudhary, and Neubig (2023);
@@
This design enables fine-grained analysis of REAL-MT robustness across both resource levels and linguistic diversity.
@@
We select three representative models to cover different model sizes and reasoning modes: two standard LLMs, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct (Team 2024), and one large reasoning model (LRM), Qwen3-8B (Team 2025). Qwen3-8B uniquely supports seamless switching between thinking mode (enabled via enable_thinking=True) and non-thinking mode (enabled via enable_thinking=False), demonstrating significantly enhanced reasoning capabilities over prior Qwen instruct models. We include this LRM to investigate whether such advanced reasoning enables models to detect and reject noisy retrieval contexts during inference.
@@
Our analysis of online idiom dictionaries and search results shows that real-world retrieval failures often arise from incomplete phrase matching or morphological variations, which frequently lead to the retrieval of either unrelated information or literal interpretations of idioms. To faithfully simulate these knowledge-level errors, we design three levels of semantic noise that form a spectrum of increasing deviation from the correct idiomatic meaning:
To balance computational cost and coverage, we randomly sample 200 instances per translation direction and use the closed-source model gemini-flash-2.0 to generate noisy contexts based on carefully designed prompt templates (see Appendix A).
@@
Across 10 language pairs, average TER($\mathcal{G}$, $\mathcal{N}_{\text{struct}}$) is 25.2, Sim($\mathcal{G}$, $\mathcal{N}_{\text{struct}}$) = 0.92, Sim($\mathcal{G}$, $\mathcal{N}_{\text{literal}}$) = 0.75, Sim($\mathcal{N}_{\text{literal}}$, $\mathcal{N}_{\text{semantic}}$) = 0.82, Sim($\mathcal{G}$, $\mathcal{N}_{\text{semantic}}$) = 0.73, and CR($\mathcal{G}$, $\mathcal{N}_{\text{opposite}}$) = 0.85, indicating that the generated noise aligns with the intended design. Full per-language results are presented in Table 1 in Appendix.
@@
Context HiEn FaEn FiEn JaEn FrEn KoEn RuEn DeEn EnFa EnDe $Avg_{C}$ $Avg_{F}$
@@
Qwen/Qwen3-8B (non-thinking mode)
@@
Qwen/Qwen3-8B (thinking mode)
@@
Table 3: Performance of LLMs on idiom translation with various context settings, C denotes Comet-22, and F denotes Fidelity.
@@
where $\text{Prompt}(y,m)$ is a prompt function that generates a text prompt for the LLM, $R$ is a random variable representing the LLM’s output. To mitigate this uncertainty and improve the accuracy of the assessment, we integrate the LLM’s output over 20 runs. The fidelity score ranges from 0 to 3, where 0 indicates completely unfaithful and 3 indicates perfectly faithful. For detailed prompt information, please refer to Appendix B.1.
@@
We design this metric to assess whether LLM utilizes the context $c$ we introduced for translation. When $c$ is noisy and the model adopts $c$, this indirectly indicates that the model lacks robustness. Given the variability in how different models translate idioms across languages and the resulting inconsistencies in fidelity quality, a single fidelity metric may not provide a clear and transparent measure of a model’s susceptibility to noise. Therefore, we introduce this metric to enhance the interpretability of how models process contextual information.
To calculate the CAR score, we first formalize the generation process of translation $y$ can be formalized as:
@@
where $y_{\text{no\_context}}$ is the translation without using context clues. We assign a score of 1 if the translation without context cues misses element $c$ while the translation using context cues successfully includes it; otherwise, a score of 0 is given. Therefore, when $c$ is noisy, the lower the CAR score, the more robust the model is. We leverage the high-performance language model gpt-4o-mini to perform this evaluation. The detailed prompt templates used for conducting evaluations are provided in Appendix B.2.
@@
Prior work shows that reference-free metrics like CometKiwi (Rei et al. 2022b) struggle to capture idiomatic meaning (Li et al. 2024). To better evaluate idiom translation, we use the reference-based COMET-22 (Rei et al. 2022a) and introduce Fidelity as a complementary metric. We validate it via human evaluation, establishing a reliable ground truth and measuring its correlation with automatic scores. We evaluate on Fi→En, Ja→En, and Fr→En, representing diverse typological families and resource levels. Translations are generated using Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, and Qwen3-8B under the No Context ($\mathcal{C}_{\text{none}}$) setting, i.e., direct translation without external knowledge. For each language pair, the first 50 instances are annotated by three linguistics experts following the same guidelines as the models (see Appendix B.1). Scores are averaged to improve reliability. As shown in Table 2, Fidelity correlates highly with human judgments across languages, demonstrating that gpt-4o-mini can serve as an effective automatic evaluator. Given this strong alignment, we adopt Fidelity as the primary metric in subsequent experiments.
To assess CAR’s reliability, we conduct a human evaluation on 200 instances, achieving 72% accuracy. Given annotation costs, we use a cost-efficient LLM-based approach. Manual analysis shows errors are mostly false 0 rather than false 1. Therefore, the high CAR values we observe under noisy contexts provide a reliable lower bound on the LLM’s reliance, strengthening the validity of our findings.
@@
We evaluate REAL-MT across six retrieval conditions: No Context ($\mathcal{C}_{\text{none}}$), which measures performance without external knowledge, and Gold Meaning ($\mathcal{G}$), which provides an upper bound using oracle idiomatic knowledge, along with four noise variants, $\mathcal{N}_{\text{literal}}$, $\mathcal{N}_{\text{semantic}}$, $\mathcal{N}_{\text{opposite}}$, and $\mathcal{N}_{\text{struct}}$, that simulate realistic retrieval failures. Comparing performance across these settings enables a systematic analysis of how noise type and severity impact translation faithfulness.
Figure 2: Context Adoption Rate (CAR) of Qwen2.5-7B-Instruct in various contexts across ten language pairs.
Knowledge-level errors severely undermine REAL-MT’s trustworthiness unlike surface-level perturbations.
As shown in Table 3, models perform best under gold meaning ($\mathcal{G}$), confirming the benefit of correct context. Performance under syntactic perturbations ($\mathcal{N}_{\text{struct}}$) is comparable to $\mathcal{G}$, indicating that the model does not reject syntactically flawed text and possesses some syntactic self-correction capabilities.. With $\mathcal{N}_{\text{literal}}$, performance slightly improves in some low-resource languages (e.g., Hi→En, Fa→En), likely due to surface-level activation of idiom-related knowledge. However, when the context is irrelevant or conveys the opposite meaning ($\mathcal{N}_{\text{opposite}}$), translations increasingly align with the corrupted context rather than the source, resulting in a performance drop. These results show that semantic noise directly undermines the trustworthiness of REAL-MT, with worse performance as semantic distortion increases.
Large reasoning models rationalize rather than reason in thinking mode.
We observe that in thinking mode, models often produce reasoning traces contradicting the noisy context, indicating awareness of inconsistency, yet still generate outputs aligned with the noisy context (see Appendix Figure 4). This behavior suggests a tendency to rationalize rather than reason, potentially due to reward hacking during training, where models prioritize contextual coherence over factual fidelity (Chen et al. 2025). Supporting this, our quantitative analysis (Table 3) shows that Qwen3-8B in non-thinking mode consistently outperforms thinking mode under noisy contexts. For instance, under $\mathcal{N}_{\text{opposite}}$, non-thinking mode achieves an average Fidelity score of $0.8$, while thinking mode also drops to $0.4$. This disconnect between reasoning and output underscores the need for truth-preserving inference in large reasoning models.
@@
Figure 2 demonstrates that medium-to-low-resource language pairs (e.g., Hi→En, Fa→En, Fi→En) show significantly higher Context Adoption Rates (CAR) compared to high-resource pairs (e.g., En→De, En→Fr, De→En) across both clean ($\mathcal{G}$) and noisy conditions ($\mathcal{N}_{\text{struct}}$, $\mathcal{N}_{\text{literal}}$, $\mathcal{N}_{\text{semantic}}$, and $\mathcal{N}_{\text{opposite}}$). The pattern generalizes to other models as detailed in Appendix C. Such heightened reliance on contextual cues suggests that low-resource languages depend more heavily on external information during translation due to less parametric idiom knowledge. Consequently, LLMs become especially vulnerable to misleading information, even when the context conveys an opposite meaning, underscoring the critical need for robustness mechanisms in low-resource machine translation settings.
@@
LLMs exhibit high Context Adoption Rates (CAR) under noisy contexts, indicating that their outputs are strongly influenced by retrieved content, as shown in Figure 2. However, CAR only measures output similarity and does not reveal why the context shapes the generation process. To uncover the underlying mechanism, we analyze attention patterns during translation, as attention is widely used to diagnose where models focus during decoding (Wiegreffe and Pinter 2019). Specifically, we compute the average cumulative attention allocated to the source idiom versus the retrieved context across all target tokens.
As shown in Figure 3, attention consistently shifts toward the contextual meaning, even when it is incorrect or adversarial. This systematic pattern indicates active integration: the model consults the context during decoding, rather than merely producing aligned output. This confirms that the model’s predictions are anchored in the provided context, even when they semantically contradict the source input.
Language Mitigation $\mathcal{C}_{\text{none}}$ $\mathcal{G}$ $\mathcal{N}_{\text{struct}}$ $\mathcal{N}_{\text{literal}}$ $\mathcal{C}_{\text{semantic}}$ $\mathcal{N}_{\text{opposite}}$
@@
Table 4: Performance of Qwen2.5-7B-Instruct after using different mitigation strategies, evaluated on three language pairs: Fr→En (high-resource), Ja→En (medium-resource), and Hi→En (low-resource).

### v2

@@ Abstract
REtrieval-Augmented LLM-based Machine Translation (REAL-MT) shows promise for knowledge-intensive tasks like idiomatic translation, but its reliability under noisy retrieval, a common challenge in real-world deployment, remains poorly understood. To address this gap, we propose a noise synthesis framework and new metrics to systematically evaluate REAL-MT’s reliability across high-, medium-, and low-resource language pairs. Using both open- and closed-sourced models, including standard LLMs and large reasoning models (LRMs), we find that models heavily rely on retrieved context, and this dependence is significantly more detrimental in low-resource language pairs, producing nonsensical translations. Although LRMs possess enhanced reasoning capabilities, they show no improvement in error correction and are even more susceptible to noise, tending to rationalize incorrect contexts. Attention analysis reveals a shift from the source idiom to noisy content, while confidence increases despite declining accuracy, indicating poor self-monitoring. To mitigate these issues, we investigate training-free and fine-tuning strategies, which improve robustness at the cost of performance in clean contexts, revealing a fundamental trade-off. Our findings highlight the limitations of current approaches, underscoring the need for self-verifying integration mechanisms.
Codehttps://github.com/ymsunny/REAL-MT-Vuln
AAAI 2026 camera-ready versionhttps://aaai.org/example/camera-ready-version
@@ Introduction
REtrieval-Augmented LLM-based Machine Translation (REAL-MT) is increasingly used to enhance translation quality for knowledge-intensive MT tasks like idiomatic translation (Li et al. 2024; Donthi et al. 2025). Although external knowledge can enhance translation performance, reliance on it is a double-edged sword: when the retrieved context contains noise such as irrelevant or misleading information, LLMs often produce nonsensical translations, as illustrated in Figure 1. Since noisy retrieval is unavoidable in real-world deployment, a critical gap remains in understanding how REAL-MT’s reliability degrades under such conditions. This limitation constitutes a key barrier to its adoption in applications where translation quality and reliability are critical. To address this, we ask: To what extent does noisy retrieval compromise REAL-MT’s trustworthiness?
@@
To systematically quantify this vulnerability, we introduce a controlled noise injection framework based on real-world retrieval failure patterns. While standard machine translation metrics like COMET (Rei et al. 2020) assess overall output quality, they fail to capture semantic fidelity in idiomatic translation and cannot distinguish whether errors stem from source misinterpretation or over-reliance on retrieved context. To address this gap, we propose two complementary metrics: Fidelity, which evaluates translation correctness with a focus on idiomatic accuracy, and Context Adoption Rate (CAR), which quantifies the extent to which models rely on external context. Together, they enable fine-grained analysis of both what the model gets wrong and why.
@@
We evaluate REAL-MT systems instantiated with both open- and closed-source models, including standard LLMs and large reasoning models (LRM), across high-, medium-, and low-resource language pairs under controlled synthesized noise. Our results show that REAL-MT is far more vulnerable to semantic-level errors, such as irrelevant or contradictory retrieval content, than to surface-level perturbations (e.g., word reordering). Moreover, the extent of performance degradation scales with the degree of semantic deviation: the more the retrieved context diverges from the intended meaning of idioms, the more severe the drop in translation fidelity. This sensitivity is especially pronounced in low-resource language pairs, as reflected in their higher CAR, which signals greater dependence on external context and leads to more severe degradation under noisy retrieval. Surprisingly, LRM, despite enhanced reasoning capabilities, show no improvement in error correction and is even more susceptible to noise, often rationalizing incorrect contexts.
Across noise conditions, CAR remains consistently high even when retrieved content contradicts the source, prompting us to examine whether this reflects active model integration or coincidental similarity. Our attention analysis confirms the former: models consistently attend to retrieved context regardless of its correctness, demonstrating active integration. This uncritical reliance is further exacerbated by poor metacognitive awareness, as confidence increases despite declining accuracy, a sign of severe miscalibration and absent self-verification.
@@
Given this uncritical reliance on retrieved context, we explore both training-free and fine-tuning strategies to improve REAL-MT robustness. While both enhance noise resistance, they incur a consistent trade-off: performance degrades under clean contexts. Fine-tuning yields better robustness overall but still fails to adjust reliance based on context quality. This persistent trade-off reveals that the current approach cannot overcome the model’s fundamental inability to self-verify, underscoring the need for self-verifying integration mechanisms that validate retrieved content before adoption.
@@
In summary, this work (1) first systematically evaluate REAL-MT robustness under noisy retrieval through a controlled noise injection framework, (2) proposes Fidelity and CAR as fine-grained metrics to quantify the impact of noisy retrieval on translation fidelity and context dependence, (3) reveal that REAL-MT overrelies on noisy context due to attention shifting from idioms to retrieved content, and becomes overconfident in wrong outputs, exposing poor calibration, and (4) evaluates training-free and fine-tuning mitigation strategies, revealing a fundamental trade-off between robustness under noise and performance in clean contexts, thereby pointing to the necessity of self-verifying integration mechanisms as a path forward.
@@
Large language models (LLMs) have revolutionized machine translation (MT), especially for low-resource languages or domains where sufficient parallel corpora are lacking (Sun et al. 2024; Zhan et al. 2021, 2024). However, when confronted with translation scenarios demanding specific background knowledge, relying solely on the LLM’s internal knowledge proves inadequate. In such instances, incorporating external knowledge to address the inherent limitations of LLMs becomes crucial (Merx et al. 2024; Chen et al. 2024; Zebaze, Sagot, and Bawden 2025). The prompt engineering capabilities of LLMs enable the integration of externally retrieved knowledge via prompting, without additional training. This makes prompt-based retrieval-augmented generation an efficient and flexible solution for MT. Recent studies by Li et al. (2024) and Donthi et al. (2024) leverage LLMs to enhance idiomatic translation by incorporating idiom-meaning pairs retrieved from offline knowledge bases directly into the prompt.
@@
Retrieval-Augmented Language Models (RALM), reliant on external retrieved content, are susceptible to compromised reliability and robustness in their generated outputs when exposed to noise, irrelevant information, or malicious data  (Zhou et al. 2024; Park and Lee 2024; Shen et al. 2024; Yang et al. 2025; Chen et al. 2025a; Wang et al. 2025). In response to the robustness challenges posed by RALMs, researchers have developed diverse approaches to enhance system robustness. Fang et al. (2024) introduces a named Retrieval-Augmented Adaptive Adversarial Training (RAAT) method to enhance the model’s ability to recognize and handle various types of noise.  Yoran et al. (2023) fine-tune the model using the parameter-efficient fine-tuning method QLoRA (Quantized Long-Range Attention) by using the synthetically generated noisy data to enhance its robustness to noisy data. Xia et al. (2025) introduce a novel end-to-end self-reasoning framework that enhances the robustness, interpretability, and traceability of RALMs. This improvement is achieved by leveraging the reasoning trajectories generated by the LLMs themselves.
@@
To analyze how resource availability affects the robustness of retrieval-augmented LLM-based machine translation (MT), we group languages into high-, medium-, and low-resource tiers following Joshi et al. (2020), based on parallel data availability. We compile a dataset of idiomatic translations across ten language pairs, including only those data sources that provide either reference translations or explicit idiom meaning annotations, ensuring each idiom has a grounded meaning annotation and facilitating subsequent controlled noise synthesis. The collection integrates multiple publicly available resources:
  • High-resource: IdiomsInCtx-MT (Stap et al. 2024) (English–German, German–English, Russian–English), the French–English, Japanese–English idiomatic translation from Liu, Chaudhary, and Neubig (2023), and the KISS dataset [^1] https://github.com/Judy-Choi/KISS-Korean-english-Idioms-in-Sentences-dataSet for Korean–English idiomatic translation;
  • Medium-resource: the Finnish–English idiomatic translation from Liu, Chaudhary, and Neubig (2023);
@@
This design enables a fine-grained analysis of REAL-MT robustness across various resource levels.
@@
We evaluate both open-source (Qwen series) and closed-source (claude-sonnet-4) models, covering standard LLMs (Qwen2.5-7B/14B-Instruct (Team 2024)) and a large reasoning model (Qwen3-8B (Team 2025)), which uniquely supports seamless switching between thinking mode (enabled via enable_thinking=True) and non-thinking mode (enabled via enable_thinking=False). This allows us to probe whether advanced reasoning capabilities improve robustness to noisy retrieval.
@@
Retrieval failures in idiomatic meaning often arise from two distinct mechanisms: incomplete phrase matching or morphological variations resulting in semantically unrelated content, while the absence of explicit idiomatic meaning online results in literal translations or low-quality paraphrases. To faithfully simulate these semantic-level errors, we design three levels of semantic noise that form a spectrum of increasing deviation from the correct idiomatic meaning:
To balance computational cost and coverage, we randomly sample 200 instances per translation direction and use gemini-flash-2.0 to generate noisy contexts based on carefully designed prompt templates in Appendix A.
@@
Across 10 language pairs, average TER($\mathcal{G}$, $\mathcal{N}_{\text{struct}}$) is 25.2, Sim($\mathcal{G}$, $\mathcal{N}_{\text{struct}}$) = 0.92, Sim($\mathcal{G}$, $\mathcal{N}_{\text{literal}}$) = 0.75, Sim($\mathcal{N}_{\text{literal}}$, $\mathcal{N}_{\text{semantic}}$) = 0.82, Sim($\mathcal{G}$, $\mathcal{N}_{\text{semantic}}$) = 0.73, and CR($\mathcal{G}$, $\mathcal{N}_{\text{opposite}}$) = 0.85, indicating that the generated noise aligns with the intended design. Full per-language results are presented in Table 5 in the Appendix.
@@
Context HiEn FaEn FiEn JaEn FrEn KoEn RuEn DeEn EnFa EnDe Avg. C Avg. F
@@
Qwen3-8B (non-thinking mode)
@@
Qwen3-8B (thinking mode)
@@
Table 3: Performance on idiom translation. Metrics include Fidelity(F), Comet-22(C), with darker colors for higher scores.
@@
where $\text{Prompt}(y,m)$ is a prompt function that generates a text prompt for the LLM, $R$ is a random variable representing the LLM’s output. We generate 20 outputs per instance and take the mode as the final score. This reduces the uncertainty in the model’s scoring as an evaluator. The fidelity score ranges from 0 to 3, where 0 indicates completely unfaithful and 3 indicates perfectly faithful. Detailed prompts, variances, and standard deviations of the evaluation scores are provided in Appendix B.1 and B.3.
@@
We design the Context Adoption Rate (CAR) metric to complement Fidelity by providing a more direct measure of a model’s susceptibility to noisy context. Fidelity alone can be ambiguous, as low scores may stem from either robust rejection of noise or poor translation quality. CAR resolves this ambiguity by directly quantifying whether the model utilizes the provided context $c$ for translation, thereby enhancing the interpretability of robustness evaluations.
To calculate the CAR score, we first formalize the generation process of translation $y$ as:
@@
where $y_{\text{no\_context}}$ is the translation without using context clues. We assign a score of 1 if the translation without context cues misses element $c$ while the translation using context cues includes it; otherwise, assign 0. Therefore, when $c$ is noisy, the lower the CAR score, the more robust the model is. We evaluate using gpt-4o-mini with prompt templates detailed in Appendix B.2.
@@
Prior work shows that reference-free metrics like CometKiwi (Rei et al. 2022b) struggle to capture idiomatic meaning (Li et al. 2024). To better evaluate idiom translation, we use COMET-22 (Rei et al. 2022a) and introduce Fidelity as a complementary metric. We validate Fidelity via human evaluation, establishing a reliable ground truth and measuring its correlation with automatic scores. We evaluate on Fi→En, Ja→En, and Fr→En, diverse in typology and resource level. Translations are generated using Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, and Qwen3-8B under the No Context ($\mathcal{C}_{\text{none}}$) setting (direct translation without external knowledge). For each language pair, the first 50 instances are annotated by three linguists following the same guidelines (see Appendix B.1) as the models. Scores are averaged for reliability. As shown in Table 2, Fidelity correlates highly with human judgments across languages, demonstrating that gpt-4o-mini is an effective automatic evaluator. We therefore adopt Fidelity as the primary metric in subsequent experiments.
To assess CAR’s reliability, we conduct a human evaluation on 200 instances, achieving 82% accuracy. Given annotation costs, we use a cost-efficient LLM-based approach (gpt-4o-mini). Manual analysis shows errors are mostly false 0 rather than false 1. Therefore, the high CAR values we observe under noisy contexts provide a reliable lower bound on LLM’s reliance, strengthening the validity of our findings. The confusion matrix between the LLM evaluator and human annotations is provided in Appendix B.4.
@@
We evaluate REAL-MT across six retrieval conditions: No Context ($\mathcal{C}_{\text{none}}$), which measures performance without external knowledge, and Gold Meaning ($\mathcal{G}$), which provides an upper bound using oracle idiomatic knowledge, along with four noise variants, Literal Translation ($\mathcal{N}_{\text{literal}}$), Semantic-Perturbed Literal Meaning ($\mathcal{N}_{\text{semantic}}$), Opposite Meaning ($\mathcal{N}_{\text{opposite}}$), and Structure-Perturbed Gold Meaning ($\mathcal{N}_{\text{struct}}$).
Figure 2: CAR scores for Qwen2.5-7B-Instruct under six retrieval conditions across ten language pairs.
Semantic-level noise severely undermines REAL-MT’s trustworthiness unlike surface-level perturbations.
As shown in Table 3, models perform best with $\mathcal{G}$ context, confirming the benefit of correct context. As shown in Figure 2 and Table 3, both the CAR and Fidelity under $\mathcal{N}_{\text{struct}}$ is comparable to that in the $\mathcal{G}$ condition, indicating that models do not reject syntactically flawed text and possess some syntactic self-correction capabilities. With $\mathcal{N}_{\text{literal}}$, performance remains largely unchanged compared to the $\mathcal{C}_{\text{none}}$ setting, indicating that idiom-relevant literal meaning has limited impact on model behavior. However, when the context is irrelevant or conveys the opposite meaning, the performance significantly drops; this sensitivity increases with the degree of semantic deviation, demonstrating the model’s non-robustness to semantic noise. Similar trends are observed in our supplementary experiments on the WMT 2023 Terminology Translation dataset using Qwen2.5-7B-Instruct model, as well as on three idiomatic translation datasets evaluated with claude-sonnet-4, supporting the generalizability of our findings to other knowledge-intensive translation tasks and model architectures. The results of supplementary experiments are provided in Appendix C.
Large reasoning model (LRM) rationalize rather than reason in thinking mode.
We observe that models often produce reasoning traces contradicting the noisy context, indicating awareness of inconsistency, yet generate outputs aligned with the noisy context (see (see Figure 8 in the Appendix D). This behavior suggests a tendency to rationalize rather than reason, potentially due to reward hacking during training, where models prioritize contextual coherence over factual fidelity (Chen et al. 2025b). This steeper performance drop occurs in thinking mode under noisy contexts. Supporting this, results in Table 3 show that Qwen3-8B in non-thinking mode consistently outperforms thinking mode, underscoring the necessity of truthful inference for LRM.
@@
Figure 2 demonstrates that medium-to-low-resource language pairs (e.g., Hi→En, Fa→En, Fi→En) show significantly higher CAR compared to high-resource pairs (e.g., En→De, En→Fr, De→En) across both clean and noisy conditions. The pattern generalizes to other models as detailed in Appendix C. Such heightened reliance on contextual cues suggests that low-resource languages depend more heavily on external information during translation due to less parametric idiom knowledge. Consequently, LLMs become especially vulnerable to misleading information, underscoring the critical need for robustness mechanisms in low-resource settings.
@@
LLMs exhibit high CAR under noisy contexts, indicating that their outputs are strongly influenced by retrieved content, as shown in Figure 2. However, CAR only measures output similarity and does not reveal why the context shapes the generation process. To uncover the underlying mechanism, we analyze attention patterns during translation, as attention is widely used to diagnose models’ focus during decoding (Wiegreffe and Pinter 2019). Specifically, we compute the average cumulative attention allocated to the source idiom versus the retrieved context across all target tokens.
As shown in Figure 3, attention consistently shifts toward the contextual meaning, even when it is adversarial. This systematic pattern indicates active integration: the model consults the context during decoding, rather than merely producing aligned output. This confirms that the model’s predictions are anchored in the provided context, even when they semantically contradict the source input.
Language Mitigation $\mathcal{C}_{\text{none}}$ $\mathcal{G}$ $\mathcal{N}_{\text{struct}}$ $\mathcal{N}_{\text{literal}}$ $\mathcal{N}_{\text{semantic}}$ $\mathcal{N}_{\text{opposite}}$
@@
Table 4: Performance of Qwen2.5-7B-Instruct after using different mitigation strategies, evaluated on three language pairs: Fr→En (high-resource), Ja→En (medium-resource), and Hi→En (low-resource). The darker colors indicate higher scores.